OpenStudy (anonymous):

how the heck do you simplify (3/16x^3)^(1/5)???

OpenStudy (anonymous):

you're not clear enough its a bit ambiguous. do you want: \[(\frac{3}{16x^3})^{\frac{1}{5}} \text{ or } (\frac{3}{16}x^3)^{\frac{1}{5}}\]

OpenStudy (anonymous):

the first one!

OpenStudy (anonymous):

well, as far as i can see there's not much you can do there to be honest

OpenStudy (anonymous):

man really? the answer is the fifth root of 6x2 all over 2x

OpenStudy (anonymous):

i mean 6x^2

OpenStudy (anonymous):

\[\sqrt[5]{\frac{ 3 }{ 16x^3 }}\] i dont really see anything else to do

OpenStudy (anonymous):

dude nickhouraney that's the original problem

OpenStudy (anonymous):

im aware lol

OpenStudy (anonymous):

hmm hold on a sec

OpenStudy (anonymous):

the answer \[\frac{\sqrt[5]{6x^2}}{2x}\] is equal to what we started with, i dont know whether i'd call it simpler.. hmm maybe

OpenStudy (anonymous):

i suppose the powers are a bit nicer

OpenStudy (anonymous):

thanks eigenschmeigen, but how did they get to that?

OpenStudy (anonymous):

Some little arrangements we have to make..

OpenStudy (anonymous):

can someone else explain? i have to go :(

OpenStudy (anonymous):

Multiply and divide by 2 and \(x^2\)..

OpenStudy (anonymous):

\[\large (\frac{3 \times 2 x^2}{16x^3 \times 2 x^2})^{\frac{1}{5}} \implies (\frac{6x^2}{2^5x^5})^{\frac{1}{5}}\]

OpenStudy (anonymous):

Now denominator will simply be taken out of the 5th root brackets : \[\large \frac{(6x^2)^{\frac{1}{5}}}{2x} \; \; Or \; \; \frac{\sqrt[5]{6x^2}}{2x}\]

OpenStudy (anonymous):

\(16 \times 2 = 32 = 2^5\)

OpenStudy (anonymous):

i'm just gonna fail this problem thanks for trying bro

OpenStudy (anonymous):

What happened ??

OpenStudy (anonymous):

Which step you are not getting ??

OpenStudy (anonymous):

i just would never think of multiplying by the x^2

OpenStudy (anonymous):

thanks though. i actually understand it kinda

OpenStudy (anonymous):

Welcome..